# Learning Loss Landscapes in Preference Optimization

### v1

@@ 2.2 Mirror Maps
Definition 2.1 ($\omega$-potential mirror map Krichene et al. (2015)).
@@
When $\phi(x)=e^{x-1}$ we recover the negative entropy mirror map, while we recover the $\ell_{2}$-norm when $\phi(x)=2x$ (refer to Appendix B). Mirror maps in this class are simple to implement in practice, where $\mathcal{A}$ is often large, as they can be parametrized by a scalar function instead of a multi-dimentional one. Additionally, the same $\omega$-potential $\phi$ can be used to generate mirror maps for different action spaces, allowing the insights obtained for one action space to easily generalize to others. An $\omega$-potential mirror map $h_{\phi}$ induces a Bregman divergence (Bregman, 1967), which is defined as
$\mathcal{D}_{h_{\phi}}(\pi(\cdot|s),\pi^{\prime}(\cdot|s)):=h_{\phi}(\pi(\cdot|s))-h_{\phi}(\pi^{\prime}(\cdot|s))-\langle\nabla h_{\phi}(\pi^{\prime}(\cdot|s)),\pi(\cdot|s)-\pi^{\prime}(\cdot|s)\rangle,$
where $\mathcal{D}_{h_{\phi}}(\pi(\cdot|s),\pi^{\prime}(\cdot|s))\geq 0$ for all $x,y\in\mathcal{Y}$. When $\phi(x)=e^{x-1}$, $\mathcal{D}_{h_{\phi}}$ is equivalent to the KL-divergence, while we recover the Euclidean distance when $\phi(x)=2x$ (refer to Appendix B). When the Bregman divergence is employed as a regularization term in optimization problems, tuning the mirror map allows us to control the geometry of the updates of the parameters to be optimized, determining when to take large or small updates based on the current value of the parameters.
3 Methodology
We develop a new framework for preference optimization based on mirror maps, which generalizes DPO and ORPO. We start by replacing the $\mathtt{KL}$-divergence penalty term in the objective in (4) with a more general Bregman divergence, that is, we aim to solve the problem
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mathcal{D},\tau\sim(\pi,P)}\left[\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}[r(s_{t},a)]-\beta D_{h}(\pi(\cdot|\tau),\pi_{\mathrm{ref}}(\cdot|\tau))\right],$ (7)
@@ 3 Methodology
We provide a proof for Theorem 3.1 in Appendix A. By plugging (8) in the maximum likelihood problem in (3), we obtain the objective:
@@
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{\mathcal{D}}\big{[}\log\sigma\big{(}\beta(\phi^{-1}(\pi(\tau_{w}))-\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{w}))-\phi^{-1}(\pi(\tau_{l}))+\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{l}))\big{)}\big{]},$ (9)
where $\mathbb{E}_{\mathcal{D}}$ is equivalent to $\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}$. When $\phi=e^{x}$, the expression in (7) recovers the DPO objective in (5). Additionally, we can modify the expression in (7) to include an SFT term in order to avoid the SFT step, in a manner similar to ORPO. Let $\psi$ be an $\omega$-potential and let $\pi_{\mathrm{ref}}$ be the uniform distribution, then the new objective we consider is
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\left[\psi(\pi(\tau_{w}))+\lambda\log\sigma\left(\phi^{-1}(\pi(\tau_{w}))-\phi^{-1}(\pi(\tau_{l}))\right)\right],$ (10)
@@
where the terms $-\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{w}))$ and $\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{l}))$ do not appear as they have cancelled out. We note that setting $\pi_{\mathrm{ref}}$ to be the uniform distribution is equivalent to replacing the Bregman divergence penalty in (7) with the mirror map $h(\pi(\cdot|\tau))$, which enforces a form of entropy regularization. When $\psi(x)=\log(x)$ and $\phi^{-1}(x)=\log(x)-\log(1-x)$, (10) recovers the ORPO objective in (6).
@@
The objectives in (9) and (10) allow us to implement a wide variety of preference optimization algorithms, while benefiting from a theoretical justification. In the following, we will show that taking into account the properties of the preference dataset when choosing $\psi$ and $\phi^{-1}$ can lead to a better performance of the trained policy.
Following the intuition developed by Jackson et al. (2024), we also allow our objective function, specifically $\psi$ and $\phi^{-1}$, to account for training progress. This is equivalent to changing the Bregman divergence penalty in (7) during training and, as we shall discuss in the following section, helps in dealing with datasets with mixed-quality data.
3.1 Learning mirror maps
@@
To search the space of PO algorithms we have defined, we employ a neural network parametrization for both $\psi$ and $\phi^{-1}$, which we optimize using evolutionary strategies.
@@
Similarly to Alfano et al. (2024), we parameterize both $\psi$ and $\phi^{-1}$ as a one layer neural network with 126 hidden units and non-negative kernels, where the activation functions are equally split among:
$x,\;(x)_{+}^{2},\;x^{3},\;(x)_{+}^{1/2},\;(x)_{+}^{1/3},\;\log((x)_{+}),\;e^{x},\;\tanh(x),\;\log(\mathop{\mathrm{clip}}(x)/(1-\mathop{\mathrm{clip}}(x))),$
@@
where $(x)_{+}=\max(x,0)$ and $\mathop{\mathrm{clip}}(x)=\max(\min(x,1),0)$. The non-negative kernels and the increasing activation functions guarantee the monotonicity of $\psi$ and $\phi^{-1}$, while the several different activation functions facilitate expressing complex functions. To ensure that we are able to recover the ORPO objective, we add $a\log(x)$ and $b\log(x)-b\log(1-x)$ to the final outputs of $\psi$ and $\phi^{-1}$, respectively, where $a,b\geq 0$. In case we want to take into account training progress, we give a second input to the neural network, i.e. $x\cdot n/N$, where $n$ is the current epoch and $N$ is the total number of epochs through the dataset. To ensure that monotonicity is preserved, we lower bound the weights associated to the second input with the negative of the respective weights associated to the first input.

### v2

@@ 2.2 Mirror Maps
Definition 2.1 ($\omega$-potential mirror map (Krichene et al., 2015)).
@@
When $\phi(x)=e^{x-1}$ we recover the negative entropy mirror map, while we recover the $\ell_{2}$-norm when $\phi(x)=2x$ (refer to Appendix D). Mirror maps in this class are simple to implement in practice, where $\mathcal{A}$ is often large, as they can be parametrized by a scalar function instead of a multi-dimentional one. Additionally, the same $\omega$-potential $\phi$ can be used to generate mirror maps for different action spaces, allowing the insights obtained for one action space to easily generalize to others. An $\omega$-potential mirror map $h_{\phi}$ induces a Bregman divergence (Bregman, 1967), which is defined as
$\mathcal{D}_{h_{\phi}}(\pi(\cdot|s),\pi^{\prime}(\cdot|s))$ $:=h_{\phi}(\pi(\cdot|s))-h_{\phi}(\pi^{\prime}(\cdot|s))$
$-\langle\nabla h_{\phi}(\pi^{\prime}(\cdot|s)),\pi(\cdot|s)-\pi^{\prime}(\cdot|s)\rangle,$
where $\mathcal{D}_{h_{\phi}}(\pi(\cdot|s),\pi^{\prime}(\cdot|s))\geq 0$ for all $x,y\in\mathcal{Y}$. When $\phi(x)=e^{x-1}$, $\mathcal{D}_{h_{\phi}}$ is equivalent to the KL-divergence, while we recover the Euclidean distance when $\phi(x)=2x$ (refer to Appendix D). When the Bregman divergence is employed as a regularization term in optimization problems, tuning the mirror map allows us to control the geometry of the updates of the parameters to be optimized, determining when to take large or small updates based on the current value of the parameters.
2.3 Evolution Strategies
OpenAI-ES  (Salimans et al., 2017) is a popular method to be able to optimize non-differentiable functions and it has been widely used to meta-learn objectives  (Lu et al., 2022; Jackson et al., 2024), as it obtains an unbiased estimate of the gradient (unlike second order gradient methods). The gradient $\nabla_{\zeta}F(\zeta)$ is estimated using:
where $\mathcal{N}(0,I_{d})$ is the multivariate normal distribution, $d$ is the number of parameters, $\widehat{F}$ is an estimate of $F$, and $\sigma>0$ is a hyperparameter regulating the variance of the perturbations.
3 Mirror Preference Optimization
We introduce Mirror Preference Optimization (MPO), a new framework for preference optimization that generalizes DPO and ORPO. We start by replacing the $\mathtt{KL}$-divergence penalty term in the objective in (4) with a Bregman divergence and aim to solve the problem
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mathcal{D},\tau\sim(\pi,P)}$ (7)
$\left[\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}r(s_{t},a)-\beta D_{h}(\pi(\cdot|\tau),\pi_{\mathrm{ref}}(\cdot|\tau))\right],$
@@ 3 Mirror Preference Optimization
We provide a proof for Theorem 3.1 in Appendix C. By plugging (11) in the maximum likelihood problem in (3), we obtain the objective:
@@
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{\mathcal{D}}\big{[}g\circ\sigma\big{(}\beta(\phi^{-1}(\pi(\tau_{w}))$ (9)
$-\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{w}))-\phi^{-1}(\pi(\tau_{l}))+\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{l}))\big{)}\big{]},$
where $\mathbb{E}_{\mathcal{D}}$ is equivalent to $\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}$.
Two-step MPO
We can use the objective in (9) to define a class of two-step PO algorithms, which consist of a preliminary SFT phase to obtain the reference policy $\pi_{\mathrm{ref}}$ and a PO phase which optimizes (9). When $\phi=e^{x}$, the (9) is equivalent to (5) and we recover the DPO algorithm.
One-step MPO
By adding an SFT term to (9) and by setting the $\pi_{\mathrm{ref}}$ be the uniform distribution, we obtain a class of one-step PO algorithms. These algorithms consists in a single phase, where we optimize the objective
$\pi^{\star}\in$ $\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big{[}\psi(\pi(\tau_{w}))$ (10)
$+\lambda g\circ\sigma\left(\phi^{-1}(\pi(\tau_{w}))-\phi^{-1}(\pi(\tau_{l}))\right)\big{]},$
@@
where $\psi$ is an $\omega$-potential. Here, the terms $-\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{w}))$ and $\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{l}))$ have canceled out due to $\pi_{\mathrm{ref}}$ being uniform. We note that setting $\pi_{\mathrm{ref}}$ to be the uniform distribution is equivalent to replacing the Bregman divergence penalty in (7) with the mirror map $h(\pi(\cdot|\tau))$, which enforces a form of entropy regularization. When $\psi(x)=\log(x)$ and $\phi^{-1}(x)=\log(x/(1-x))$, (10) recovers the ORPO objective in (6).
@@
The objectives in (9) and (10) allow us to implement a wide variety of preference optimization algorithms, while benefiting from a theoretical justification. In the following, we will show that it is possible to parametrize and optimize $g$, $\psi$, and $\phi^{-1}$ to obtain new algorithms that outperform current baselines.
3.1 Meta Learning PO objectives
@@
To search the space of PO algorithms we have defined, we employ a neural network parametrization for $g$, $\psi$, and $\phi^{-1}$, which we optimize using evolutionary strategies (Salimans et al., 2017).
@@
Similarly to (Alfano et al., 2024), we parameterize $g$, $\psi$ and $\phi^{-1}$ as a one layer neural network with 126 hidden units and non-negative kernels, where the activation functions are equally split among:
$x,\;(x)_{+}^{2},\;x^{3},\;(x)_{+}^{1/2},\;(x)_{+}^{1/3},\;\log((x)_{+}),$
$e^{x},\;\tanh(x),\;\log(\mathop{\mathrm{clip}}(x)/(1-\mathop{\mathrm{clip}}(x))),$
@@
where $(x)_{+}=\max(x,0)$ and $\mathop{\mathrm{clip}}(x)=\max(\min(x,1),0)$. The non-negative kernels and the increasing activation functions guarantee the monotonicity of $g$, $\psi$, and $\phi^{-1}$, while the several different activation functions facilitate expressing complex functions. To ensure that we are able to recover the DPO and ORPO objectives, we add $a\log(x)$, $b\log(x)$ and $c\log(x/(1-x))$ to the final outputs of $g$, $\psi$ and $\phi^{-1}$, respectively, where $a,b,c\geq 0$.